OpenStudy (anonymous):

If f(x) = 5x2 − 4 and g(x) = x − 2, what is (f o g)(-3)?

OpenStudy (anonymous):

You need to compose the functions f and g. That is,\[fog(x)=f(g(x))=5[g(x)]^2-4=5(x-2)^2-4\]

OpenStudy (anonymous):

You then need to evaluate at x=-3:\[fog(-3)=5(-3-2)^2-4=5(25) - 4 =121\]